# A Review of Hydrogen-Enabled Resilience Enhancement for Multi-Energy Systems

### v2

@@ VII-B Multi-stage operation strategy
Figure 15: The framework of resilient scheduling for electricity-hydrogen distribution networks
Summary: For easy understanding of existing works related to resilient operation for HMESs, we summarize their characteristics in Table IX and Table X, and list hydrogen-related measures in each response stage in Table XI. It can be observed that hydrogen plays an important role in every response stage, such as pre-charging HESS in the preventive response stage, hydrogen for fuel support via natural gas pipeline in the emergency response stage, and scheduling MHERs in the restoration response stage. Moreover, existing works mainly focus on single-stage or two-stage operation methods. Few of them consider the coordination operation among three kinds of response stages. In addition, hydrogen pipeline-related operational measures in emergency response and restoration response could be observed. The reason is that hydrogen pipelines can serve as large-scale energy storage solutions and play a key role in resilience enhancement. As mentioned in existing work[119],[124], the pressure of a hydrogen pipeline can be adjusted dynamically by changing the quantity of hydrogen within it. By exploiting such elasticity, many resilience enhancement measures related to hydrogen pipelines could be adopted. For example, MESs can temporarily store hydrogen in the pipeline when there are extra renewable energy but limited hydrogen tank space. When extreme event happens, the hydrogen pipeline can release hydrogen to drive fuel cell for energy supply. Such linepack utilization is helpful for enhancing resilience of MESs[125]. Moreover, during extreme events (such as main grid outages or natural gas supply interruptions), hydrogen pipelines can be directionally dispatched to recover hydrogen supply to critical loads such as hospitals and communication base stations. Finally, most contingencies in hydrogen-enabled resilience enhancement for MESs in existing works are related to main grid power outages, component outages, renewable generation reduction, and load surge, cyber-attack related contingencies should be paid more attention, which are also very common in cyber-physical energy systems[126, 127, 128, 129].
@@ VIII-A Comprehensive Design of HMES Resilience Metrics
VIII-B Temporally-correlated Event-targeted Scenario Generation for Planning
In existing works on planning methods for the resilience enhancement of HMESs, uncertainties related to system parameters (e.g., distributed generation output and versatile demands) are seldom considered, or just several representative days are considered. In fact, the contingency duration incurred by an extreme event (e.g., extreme snow and freezing ice disasters) may last for several days[10][131]. Thus, the representative day is not long enough to represent the contingency uncertainty. Moreover, when taking system parameters into consideration, their correlations should be utilized to generate more representative and practical scenarios. For example, in some places, when heat demand is needed in winter, the corresponding cold demand may be not required. Similarly, when cold demand is needed in summer, the corresponding heat demand may be negligible[132]. In addition, extreme events generally have spatial features and their uncertainties should also be considered for scenario generation[47]. In summary, how to generate temporally-correlated and event-targeted scenarios for planning method design requires further investigation. A possible solution for generating temporally-correlated scenarios is to generate normalized multi-parameter vector scenarios according to the duration of extreme events and condense them using scenario reduction methods. Finally, the condensed normalized multi-parameter vectors are recovered using denormalization operations.
@@ VIII-D Multi-network Multi-timescale Coordination Planning and Operation
Most existing works mainly focus on enhancing the resilience of HMESs related to electricity demand and hydrogen demand. However, two aspects are neglected. Firstly, multi-timescale feature existing in electric-hydrogen systems is not considered. To be specific, slow dynamics of hydrogen systems and the fast response of electric systems are coupled with each other. For instance, hydrogen pipelines exhibit response delays and they can not respond instantly to changes in flow or pressure, i.e., the pipeline system does not reach a new equilibrium immediately after a disturbance, and it takes time for changes in input or demand to affect the system’s operation. Such delay may span from minutes to hours. In addition to slow dynamics, hydrogen pipeline can be used to store hydrogen temporarily due to the compressibility of hydrogen, which can help to absorb fluctuations in renewable energy generation or load demand without the immediate need for additional storage tanks. Compared with hydrogen tank/pipelines, large-scale hydrogen energy storage systems (e.g., salt-cavern hydrogen storage) have slower dynamics, which are used as SHSs to adjust the weekly/monthly source-load imbalance[54]. In contrast, to capture the minute-level fluctuations of renewable energy and load demand, electricity system could be scheduled at minute timescale. Thus, multi-timescale modeling should be considered in the planning and operation of hydrogen-enabled MESs. Secondly, the same attention should be paid to providing heat demand and gas demand with high service quality under extreme events, since such demands may exceed 50% of total energy demand in a country[108], e.g., almost half of the energy demand in the U.K. is attributed to heat demand. When HMESs with versatile energy demands are considered, more strong couplings of power/gas/heat networks would be incurred. To manage such systems of systems efficiently, multi-network coordination is required. since enhancing the HMES resilience requires the use of many MHERs (e.g., HTTs, HFCVs, and HFCBs) and other mobile electric energy resources (e.g., mobile wind turbines and electric vehicles), four-network (i.e., power/gas/heat/transportation networks) coordination in planning and operation deserves further investigation. To support multi-timescale coordination operation, a possible solution is to utilize hierarchical MPC or hierarchical deep reinforcement learning for multi-timescale coordination operation, e.g., deciding the quantity of MHERs at day timescale, scheduling MHERs in emergency response at the hourly timescale, and scheduling electric equipment at the 5-minute timescale.
@@ VIII-F LLM-assisted Whole-Process Resilience Enhancement for HMESs
LLMs (e.g., GPT-4) have demonstrated remarkable capabilities in understanding, synthesizing, and reasoning across vast domains of unstructured and structured information[149, 150]. Their contextual awareness, semantic reasoning, and adaptability position them as promising tools for supporting the whole-process resilience enhancement of HMESs. Firstly, the current resilience metrics failed to capture the hydrogen-related attributes, e.g., hydrogen availability, tank/pipeline pressure recovery, and sector-coupled response. In contrast, LLMs can extract, analyze, and generate multi-dimensional, cross-sectoral resilience indicators by mining technical literature, policy documents, and incident reports, providing a foundation for standardizing hydrogen-specific evaluation frameworks. Secondly, contingency modeling under extreme events in HMESs is hindered by the complexity of hydrogen-related subsystems, while LLMs can contribute to contingency modeling by constructing structured fault trees and scenario libraries based on natural language descriptions, enabling the automatic generation of probabilistic or rule-based failure scenarios under diverse stress conditions. Thirdly, from a planning enhancement perspective, LLMs can support scenario generation and strategic siting of hydrogen assets by reasoning over multi-objective (e.g., operation cost, carbon emission, and resilience) trade-offs. Finally, from the operation enhancement perspective, LLMs offer potential in assisting the dynamic dispatch and routing of MHERs since they can interpret complex, multi-modal inputs (ranging from sensor data to emergency protocols) and contribute to human-in-loop decision.
@@ IX Conclusions
This article presents the first comprehensive review of hydrogen-enabled resilience enhancement for MESs from the perspective of planning and operation. Firstly, advantages and challenges of adopting hydrogen to enhance resilience of MESs are introduced. Then, it provides a comprehensive resilience enhancement framework for HMESs. Under the proposed framework, the widely used resilience metrics and event-oriented contingency models are summarized. Then, it classifies the hydrogen-enabled planning measures for resilience enhancement of HMESs according to the type of hydrogen-related facilities and summarizes the typical planning problem formulation frameworks. Next, it categorizes the hydrogen-enabled operation measures for the resilience enhancement of HMESs according to three types of operation response stages and summarizes the role of hydrogen in each response stage. Finally, several research gaps and future directions are identified, i.e., (1) Comprehensive HMES resilience metrics with the consideration of hydrogen-specific attributes (e.g., hydrogen system availability ratio and pressure recovery time), load loss, the number of customers affected, and carbon resilience should be designed; (2) Generating temporally-correlated event-targeted scenario generation for planning based on normalized multi-parameter vector scenarios and scenario reduction methods; (3) Achieving multi-type temporal-spatial contingency modeling under compound extreme events based on spatially-aware dynamic Bayesian networks and Petri-net; (4) Implementing multi-network multi-timescale coordination planning and operation based on hierarchical MPC or hierarchical deep reinforcement learning; (5) Designing advanced artificial intelligence-based multi-objective optimization methods for the planning and operation of low-carbon and resilient HMES; (6) Adopting LLM-assisted whole-process resilience enhancement for HMESs in aspects of proper resilience metrics, event-oriented contingency modeling, scenario generation and smart decision in planning process, and MHER scheduling and routing in operation process. While the above-mentioned technical strategies are essential for enhancing the resilience of HMESs, business and market models also play a crucial role. Representative models include hydrogen-as-a-service, priority hydrogen supply contracts, and shared hydrogen pipeline infrastructure. For instance, by leasing hydrogen fuel cell trucks under a hydrogen-as-a-service model, MESs can gain black-start capabilities and rapidly restore critical loads during unexpected power outages without requiring large upfront investments. Likewise, priority hydrogen contracts can ensure uninterrupted supply to critical infrastructure (e.g., hospitals or water treatment plants) during extreme events. Moreover, due to the high capital cost of hydrogen pipeline infrastructure, shared investment models among multiple MES operators can promote cost-effective deployment while improving cross-system resilience.

### v3

@@ VII-B Multi-stage operation strategy
Figure 15: The framework of resilient scheduling for electricity-hydrogen distribution networks[77]
Summary: For easy understanding of existing works related to resilient operation for HMESs, we summarize their characteristics in Table XI and Table XII, and list hydrogen-related measures in each response stage in Table XIII. It can be observed that hydrogen plays an important role in every response stage, such as pre-charging HESS in the preventive response stage, hydrogen for fuel support via natural gas pipeline in the emergency response stage, and scheduling MHERs in the restoration response stage. Moreover, existing works mainly focus on single-stage or two-stage operation methods. Few of them consider the coordination operation among three kinds of response stages. In addition, hydrogen pipeline-related operational measures in emergency response and restoration response could be observed. The reason is that hydrogen pipelines can serve as large-scale energy storage solutions and play a key role in resilience enhancement. As mentioned in existing work[136],[140], the pressure of a hydrogen pipeline can be adjusted dynamically by changing the quantity of hydrogen within it. By exploiting such elasticity, many resilience enhancement measures related to hydrogen pipelines could be adopted. For example, MESs can temporarily store hydrogen in the pipeline when there is surplus renewable energy but limited hydrogen tank space. When an extreme event happens, the hydrogen pipeline can release hydrogen to drive fuel cells for energy supply. Such linepack utilization is helpful in enhancing resilience of MESs[141]. Moreover, during extreme events (such as main grid outages or natural gas supply interruptions), hydrogen pipelines can be directionally dispatched to recover hydrogen supply to critical loads such as hospitals and communication base stations. Finally, most contingencies in hydrogen-enabled resilience enhancement for MESs in existing works are related to main grid power outages, component outages, renewable generation reduction, and load surge, cyber-attack related contingencies should be paid more attention, which are also very common in cyber-physical energy systems[142, 143, 144, 145].
@@ VIII-A Comprehensive Design of HMES Resilience Metrics
VIII-B Extreme-event Scenario Generation for HMES Resilience Planning
In existing studies on HMES resilience planning, only a limited number of works have explicitly considered extreme-event uncertainty. Among them, extreme-event scenario generation has mainly been conducted using generative adversarial networks (GANs) [47] and climate- and weather-model-driven methods [105]. GAN-based approaches are attractive because they can learn high-dimensional spatio-temporal patterns from data and generate large numbers of synthetic extreme scenarios, which is particularly useful for HMESs where prolonged renewable deficits may coincide with heating or cooling stress. Climate- and weather-model-driven methods, by contrast, can provide physically grounded realizations of persistent meteorological extremes, such as prolonged heatwaves, droughts, and renewable-energy droughts, which are often insufficiently represented in historical observations. Despite these advantages, both approaches have important limitations, as summarized in Table X. In particular, GAN-based methods are typically data-hungry and may generate physically implausible scenarios in the absence of sufficient domain constraints. Climate- and weather-model-driven methods, on the other hand, suffer from high computational burden, the need for downscaling and bias correction, the difficulty of translating raw climate outputs into planning-ready scenarios for HMESs, and structural uncertainty arising from imperfect or simplified model formulations. More importantly, these limitations are not confined to the above two categories, but are also relevant to other extreme-event scenario generation methods listed in Table X. Among them, one of the most critical unresolved issues is extreme-data scarcity. By definition, severe extreme events are rare, while compound extremes are even less frequently observed. As a result, the available data record is often too short to support robust characterization of long-duration, spatially correlated, and concurrent extreme events. This challenge is particularly pronounced for HMESs, because the resilience value of hydrogen is expected to emerge precisely under low-frequency but high-impact events, whereas such conditions are poorly represented in ordinary operational datasets. Therefore, addressing extreme-data scarcity is a key prerequisite for more credible resilience-oriented planning. Several promising directions may help address this challenge. First, extreme-value-theory-assisted deep generative models offer a valuable hybrid framework, in which deep generative models capture high-dimensional dependence structures while extreme value theory improves the representation and extrapolation of tail behavior [125]. Second, physics-informed generative models provide a way to enhance the physical plausibility of synthetic extreme scenarios by embedding conservation laws, known meteorological drivers, or other domain constraints into the generation process [125]. Third, response-aware scenario generation methods can identify representative climate scenarios according to system response, thereby reducing scenario complexity while preserving the most critical stress patterns for resilience-oriented planning [132]. These directions suggest that future extreme-event scenario generation for HMES resilience planning should move beyond purely data-fitting approaches toward hybrid frameworks that are tail-aware, physically consistent, and decision-relevant.
@@ VIII-D Multi-network Multi-timescale Coordination Planning and Operation
Most existing works mainly focus on enhancing the resilience of HMESs related to electricity demand and hydrogen demand. However, two aspects are neglected. Firstly, multi-timescale feature existing in electric-hydrogen systems is not considered. To be specific, slow dynamics of hydrogen systems and the fast response of electric systems are coupled with each other. For instance, hydrogen pipelines exhibit response delays and they can not respond instantly to changes in flow or pressure, i.e., the pipeline system does not reach a new equilibrium immediately after a disturbance, and it takes time for changes in input or demand to affect the system’s operation. Such delay may span from minutes to hours. In addition to slow dynamics, hydrogen pipelines can be used to store hydrogen temporarily due to the compressibility of hydrogen, which can help to absorb fluctuations in renewable energy generation or load demand without the immediate need for additional storage tanks. Compared with hydrogen tanks/pipelines, large-scale hydrogen energy storage systems (e.g., salt-cavern hydrogen storage) have slower dynamics, which are used as SHSs to adjust the weekly/monthly source-load imbalance[54]. In contrast, to capture the minute-level fluctuations of renewable energy and load demand, electricity system could be scheduled on a minute timescale. Thus, multi-timescale modeling should be considered in the planning and operation of hydrogen-enabled MESs. Secondly, the same attention should be paid to providing heat demand and gas demand with high service quality under extreme events, since such demands may exceed 50% of total energy demand in a country[109], e.g., almost half of the energy demand in the U.K. is attributed to heat demand. When HMESs with versatile energy demands are considered, more strong couplings of power/gas/heat networks would arise. To manage such systems of systems efficiently, multi-network coordination is required. since enhancing the HMES resilience requires the use of many MHERs (e.g., HTTs, HFCVs, and HFCBs) and other mobile electric energy resources (e.g., mobile wind turbines and electric vehicles), four-network (i.e., power/gas/heat/transportation networks) coordination in planning and operation deserves further investigation. To support multi-timescale coordination operation, a possible solution is to utilize hierarchical MPC or hierarchical deep reinforcement learning for multi-timescale coordination operation, e.g., deciding the quantity of MHERs at day timescale, scheduling MHERs in emergency response at the hourly timescale, and scheduling electric equipment at the 5-minute timescale.
@@ VIII-F LLM-assisted Whole-Process Resilience Enhancement for HMESs
LLMs (e.g., GPT-series models) have demonstrated remarkable capabilities in understanding, synthesizing, and reasoning across vast domains of unstructured and structured information[164, 165]. Their contextual awareness, semantic reasoning, and adaptability position them as promising tools for supporting the whole-process resilience enhancement of HMESs. Firstly, the current resilience metrics failed to capture the hydrogen-related attributes, e.g., hydrogen availability, tank/pipeline pressure recovery, and sector-coupled response. In contrast, LLMs can help extract and organize hydrogen-specific resilience attributes from technical literature, safety standards, policy documents, and incident reports, thereby supporting the development of more comprehensive resilience metrics for HMESs. Secondly, contingency modeling under extreme events in HMESs is hindered by the complexity of hydrogen-related subsystems. For example, during a typhoon event, transmission outages may coincide with road blockages, temporary natural gas shortages, and constrained hydrogen delivery, creating cascading impacts that are difficult to model using only predefined structured data. In this context, LLMs can assist contingency modeling by converting natural-language descriptions from operation manuals, fault logs, and emergency response documents into structured fault trees, event chains, and scenario libraries, which can improve the efficiency of constructing probabilistic or rule-based contingency models under diverse stress conditions. Thirdly, from a planning perspective, LLMs can support semantically grounded extreme-event scenario generation since their training are conducted based on massive textual corpora[125]. In addition, LLMs can support strategic siting of hydrogen-related assets by transforming multi-source textual knowledge into candidate rules, constraints, and evaluation criteria. For instance, an LLM may infer that a location close to wind power resources is economically attractive for electrolysis, but unsuitable for emergency hydrogen supply if it is far from hospitals, communication hubs, or major transport corridors. Finally, from an operational perspective, LLMs can support the dynamic dispatch of HMESs by integrating heterogeneous real-time information (e.g., sensor readings, traffic updates, weather alerts, and emergency protocols) and assisting decision-making on hydrogen allocation, mobile resource routing, and recovery prioritization. Therefore, rather than replacing physics-based models or optimization tools, LLMs can serve as an intelligent interface that connects fragmented knowledge, planning rules, and real-time operational information, thereby enhancing the resilience of HMESs across assessment, modeling, planning, and operation.
@@ IX Conclusions
This article presents the first comprehensive review of hydrogen-enabled resilience enhancement for MESs from the perspective of planning and operation. Firstly, advantages and challenges of adopting hydrogen to enhance resilience of MESs are introduced. Then, it provides a comprehensive resilience enhancement framework for HMESs. Under the proposed framework, the widely used resilience metrics and event-oriented contingency models are summarized. Then, it classifies the hydrogen-enabled planning measures for resilience enhancement of HMESs according to the type of hydrogen-related facilities and summarizes the classifications of uncertainty processing methods and scenario generation methods, as well as planning problem formulation frameworks. Next, it categorizes the hydrogen-enabled operation measures for the resilience enhancement of HMESs according to three types of operation response stages and summarizes the role of hydrogen in each response stage. Finally, several research gaps and future directions are identified, i.e., (1) Comprehensive HMES resilience metrics with the consideration of hydrogen-specific attributes (e.g., hydrogen system availability ratio and pressure recovery time), load loss, the number of customers affected, and carbon resilience should be designed; (2) Generating tail-extrapolation-aware, physical-informed, response-relevant extreme-event scenarios for HMES resilient planning based on extreme value theory-assisted deep generative models, physics-informed generative models, and response-aware scenario generation methods; (3) Achieving multi-type temporal-spatial contingency modeling under compound extreme events based on spatially-aware dynamic Bayesian networks and Petri-net; (4) Implementing multi-network multi-timescale coordination planning and operation based on hierarchical MPC or hierarchical deep reinforcement learning; (5) Designing advanced artificial intelligence-based multi-objective optimization methods for the planning and operation of low-carbon and resilient HMES; (6) Adopting LLM-assisted whole-process resilience enhancement for HMESs in aspects of proper resilience metrics, event-oriented contingency modeling, scenario generation and smart decision in planning process, and MHER scheduling and routing in operation process. While the above-mentioned technical strategies are essential for enhancing the resilience of HMESs, business and market models also play a crucial role. Representative models include hydrogen-as-a-service, priority hydrogen supply contracts, and shared hydrogen pipeline infrastructure. For instance, by leasing hydrogen fuel cell trucks under a hydrogen-as-a-service model, MESs can gain black-start capabilities and rapidly restore critical loads during unexpected power outages without requiring large upfront investments. Likewise, priority hydrogen contracts can ensure uninterrupted supply to critical infrastructure (e.g., hospitals or water treatment plants) during extreme events. Moreover, due to the high capital cost of hydrogen pipeline infrastructure, shared investment models among multiple MES operators can promote cost-effective deployment while improving cross-system resilience.